Suppose 2x−3y = −15 and 6x− 8y+z = 64. Find the whole number value of y+ z.

Grade:9

1 Answers

Arun
6 years ago
Dear Shrey
 
From 1st equation
2x = 3y -15
So
6x = 9y - 45
Put this is equation 2nd
9y - 45 - 8y +z = 64
y +z = 64 +45 = 109
y + z = 109
 
Regards
Arun (askIITians forum expert)